In the binomial expansion of (a - b)n, n = 5, the sum of 5th and 6th terms is 0, then a / b equals

Grade:12

4 Answers

bharat bajaj IIT Delhi
9 years ago
t5+ t6= 0
nC4an - 4(- b)4+nC5an - 5(- b)5= 0
Solving above equation, we get a / b = (n - 4) / 5
Thanks
Bharat Bajaj
askiitians faculty
Ujwal hirishi
6 years ago
by using easy method
tr+1/tr = n-r+1/r
=>n-5+1/5=n-4/5 for r =5
it is a easy method which can be used to solve any iit problem
Vaishali Senad
5 years ago
A/C to question ;
t5 + t6 = 0 ——— (1)
Since, t(r+1)= nCr. a^(n-r). b^r
Therefore, from (1) ;
nC4. a^(n-4). (-b)^4 + nC5. a^(n-5). (-b)^5=0
 
nC4. a^(n-4). b^4=
Vaishali Senad
5 years ago
Sorry for incomplete answer
nC4. a^(n-4). b^4= nC5. a^(n-5). b^5
[a^(n-4). b^4] / [a^(n-5). b^5]= nC5/nC4
a^(n-4-n+5). b^(4-5)= nC5/nC4
So, a/b= nC5/ nC4
a/b = n!/(n-5)!4! x (n-4)!4!/n!
Therefore,  " a/b= (n-4)/5. "